The following table contains information on matched sample values upon which the differences

calculated between the two samples are assumed to be normally distributed.

Number 1 2 3 4 5 6 7 8

Sample 1 18 12 21 22 16 14 17 18

Sample 2 21 11 23 20 20 17 17 22

The E–Tutor wants to test whether the mean difference differ from zero.

Which one of the following statements is incorrect?

1. The hypothesis are: H0 : D D 0 vs H1 : D 6D 0:

2. The sample mean is .


1
Expert's answer
2021-09-02T15:43:46-0400

Statement 1

"H_o: \mu_0=0\\
H_1: \mu_0=0\\"


Statement 2

"\bar{d}= \frac{\sum d_i}{n}= -1.625"

"s.d = \sqrt{\frac{\sum d_i^2-(d_i)^2}{n-1}}= 2.3261"


Statement 3

"s_e= \frac{s_d}{\sqrt{n}}= \frac{2.3261}{\sqrt{8}}=0.8224"


Statement 4

"t= \frac{\bar{d}}{s_e}= \frac{-1.625}{0.8224}=-1.9759"


Statement 5

Reject if t < -t7

t< -3.499

Hence the fifth statement is false